# TrustLLM: Smelling and Reasoning Smart Contract Vulnerability

### v1

@@ Abstract
In this paper, we propose TrustLLM, a general framework that combines fine-tuning and LLM-based agents for intuitive smart contract auditing with justifications. Specifically, TrustLLM is inspired by the observation that expert human auditors first perceive what could be wrong and then perform a detailed analysis of the code to identify the cause. As such, TrustLLM employs a two-stage fine-tuning approach: it first tunes a Detector model to make decisions and then tunes a Reasoner model to generate causes of vulnerabilities. However, fine-tuning alone faces challenges in accurately identifying the optimal cause of a vulnerability. Therefore, we introduce two LLM-based agents, the Ranker and Critic, to iteratively select and debate the most suitable cause of vulnerability based on the output of the fine-tuned Reasoner model. To evaluate TrustLLM, we collected a balanced dataset with 1,734 positive and 1,810 negative samples to fine-tune TrustLLM. We then compared it with traditional fine-tuned models (CodeBERT, GraphCodeBERT, CodeT5, and UnixCoder) as well as prompt learning-based LLMs (GPT4, GPT-3.5, and CodeLlama-13b/34b). On a dataset of 263 real smart contract vulnerabilities, TrustLLM achieves an F1 score of 91.21% and an accuracy of 91.11%. The causes generated by TrustLLM achieved a consistency of about 38% compared to the ground truth causes.
@@ I Introduction
Fine-tuning [17, 18] could be a promising approach to embed Solidity-specific vulnerability data into the model itself, compared to RAG [19], and thus address the problem mentioned above. In particular, by fine-tuning an LLM with vulnerable and non-vulnerable code, it could effectively perceive whether a new piece of code is vulnerable or not. According to insights from a million-dollar-earning hacker mentioned in the prologue, such intuition is quite important for vulnerability auditing. As such, instead of fine-tuning a single model to generate both vulnerability decisions (i.e., Yes or No) and the causes of vulnerabilities (i.e., the type or reason) simultaneously, we propose a novel two-stage fine-tuning approach. This approach first tunes a Detector model to make decisions only, and then tunes a Reasoner model to generate the causes of vulnerabilities. In this way, the fine-tuned LLMs could could mimic human hackers by first making intuitive judgments and then performing follow-up analysis of the code to identify the reasons for vulnerabilities.
@@
We implement this “perception-then-analysis” fune-tuning into a general framework called TrustLLM for intuitive smart contract auditing. In this implementation, TrustLLM allows Detector to make multiple intuitive judgments, each representing one perception. To achieve this, TrustLLM generates multiple variant prompts for the same vulnerability label to tune Detector and similarly employs multiple variant prompts for the same vulnerability reason to tune Reasoner. While it is possible to determine the optimal decision based on majority voting, fine-tuning alone cannot identify the optimal cause for a vulnerability during the inference phase. To address this new problem, we introduce the concept of LLM-based agents to the paradigm of fine-tuning in TrustLLM. Specifically, we introduce two dedicated LLM-based agents, the Ranker and Critic agents, to iteratively select and debate the most appropriate cause of vulnerability based on the output of the fine-tuned Reasoner model.
@@
To obtain high-quality data for training and testing TrustLLM, we propose leveraging reputable auditing reports to collect positive samples and employing our own data enhancement method to derive negative samples. Eventually, we collected a balanced dataset consisting of 1,734 positive samples, i.e., vulnerable functions with reasons from 263 smart contract auditing reports, and 1,810 negative samples, i.e., non-vulnerable benign code. We then compared TrustLLM with traditional full-model fine-tuning methods, including CodeBERT, GraphCodeBERT, CodeT5, and UnixCoder, as well as with prompt learning-based LLMs, such as GPT-4/GPT-3.5 and CodeLlama-13b/34b. Our experimental results show that TrustLLM achieved an F1 score of 91.21%, significantly outperforming prompt learning-based LLMs (which are in the range of 60%+) and also notably beating other fine-tuned models (which are in the range of 80%+) that used the same training data as ours. Furthermore, in terms of alignment with ground-truth explanations, TrustLLM’s output is clearly superior to that of other models, reaching a consistency rate of 37.99%. In contrast, the second-ranked GPT-4 achieves only 24%.
@@
Besides the evaluation results, we also conducted three ablation studies to further justify TrustLLM’s two-stage fine-tuning and majority voting strategies, as well as to measure the impact of additional call graph information on the model’s performance. We summarize the key findings as follows:
@@
  • TrustLLM’s two-stage approach achieved better detection performance than the integration model, which outputs labels and reasons simultaneously. We also experimentally confirmed that the model struggles to focus on the labels when required to output both types of information.
@@
Roadmap. The rest of this paper is organized as follows. We first introduce the relevant background in §II, followed by the design of TrustLLM in §III. We then present our experimental setup and the results in §IV. After that, we discuss related work and the limitations in §V and §VI, respectively. Finally, §VII concludes this paper.
@@ II-C Smart Contracts and Their Vulnerabilities
Figure 1: An overview of TrustLLM, featuring its four roles: Detector, Reasoner, Ranker, and Critic.
III Design of TrustLLM
As motivated in §I, TrustLLM employs a novel two-stage fine-tuning approach and combines it with LLM-based agents for intuitive smart contract auditing with justifications. As shown in Fig. 1, TrustLLM has the following four roles:
@@ III Design of TrustLLM
  • Reasoner takes Detector’s initial vulnerability perception to further investigate problematic code locations. By connecting Detector’s output with Reasoner’s reasoning during both training and inference, TrustLLM achieves two-stage fine-tuning.
  • To identify the optimal cause of a vulnerability during the inference phase, we further introduce the concept of LLM-based agents into the fine-tuning paradigm in TrustLLM. Specifically, Ranker evaluates the reasons for each potential vulnerability, selecting a top explanation, while Critic further assesses Ranker’s output to debate and determine the most appropriate cause of the vulnerability.
Challenges. While TrustLLM’s four roles in Fig. 1 are intuitive, training and coordinating them well for effective smart contract auditing with reasonable justifications is difficult. More specifically, we encountered the following four challenges during the design and implementation of TrustLLM:
@@
How to collect and derive high-quality training data? For a fine-tuned model like TrustLLM, obtaining high-quality training data is always crucial. We propose leveraging reputable auditing reports to collect positive samples and employing our own data enhancement method to derive negative samples. Since this part is independent of TrustLLM’s design, we defer its presentation to the end of this section in §III-D.
@@
How to make effective vulnerability judgements? While fine-tuning a model with vulnerable and non-vulnerable code is straightforward, tuning it to be effective with limited data presents a challenge. We make an effort towards addressing this problem in §III-A by opting to use multiple prompts for fine-tuning rather than a single prompt. The advantages of this approach are twofold: (i) it enriches the training dataset by increasing the volume of data, and (ii) it diminishes the bias associated with a single prompt, thereby enhancing the reliability of the results [48]. Optimal vulnerability perception could thus be achieved through majority voting.
@@
How to effectively connect Detector’s vulnerability sensing with Reasoner’s vulnerability reasoning? The fine-tuning of TrustLLM is unique because it employs a two-stage fine-tuning approach with the Detector and Reasoner models. Therefore, how to effectively connect these two models becomes a new issue not encountered in traditional fine-tuning. We present this aspect of TrustLLM’s design in §III-B.
@@
How to obtain the optimal vulnerability cause from Reasoner’s output? Since Reasoner also employs multi-prompt fine-tuning, it is necessary to identify the optimal cause of vulnerability among the multiple causes output by Reasoner. We introduce two LLM-based agents, namely the Ranker and Critic components, in §III-C, to iteratively select and debate the most appropriate cause of vulnerability.
An Example of Workflow. To wrap up, Fig. 1 also illustrates an example of TrustLLM’s workflow. Initially, Detector perceives code vulnerabilities using five different inference paths (prompts). The perceived results are then subjected to majority voting to determine a consensus label. Based on the voting result, Reasoner interprets this outcome according to different inference paths, resulting in ten answers (each considering the context of the code location or not). Next, Ranker selects Reason 1 with a confidence score 9/10 and explains this choice. Critic challenges this choice and advises Ranker to re-evaluate. Taking Critic’s feedback into account, Ranker re-ranks the ten reasons and selects Reason 3 with a confidence score of 10/10. Critic reviews Ranker’s choice again and agrees with this decision. The loop is completed, and the final reason is returned to the user.
@@ III-A Using Multi-prompt Tuning and Majority Voting for Effective Vulnerability Judgements in the Detector
Detector is a fine-tuned expert model responsible for assessing whether code poses any risk. It mimics human intuitive judgment upon seeing a piece of code, assessing whether there are any issues. We employed LoRA [39] to fine-tune CodeLlama-13b [49] in the instruction manner [50] based on a high quality of dataset. During training, for the same input code, we wrap it with multiple prompts. These prompts, with different instruction formats, represent the different inference paths, as illustrated in Fig. 1. In the inference phase of Detector, based on the output results of each prompt, we adopt a majority voting approach to determine the input label and use the voting ratio as the confidence score. Based on Detector’s majority voting result, Reasoner in §III-B then generates different reasons according to different inference paths.
Below is an instruction that describes a classification task. [Task Description] ### Instruction:
### Input: : “‘Solidiy {code} ”’ ### Response:
@@ III-B Connecting Detector for Reasoner’s Tuning & Inference
Reasoner is an expert model responsible for reasoning about and explaining code vulnerabilities. It interprets the majority voting result of the Detector, generating multiple alternative explanations. During the Reasoner’s LoRA fine-tuning process, our inputs include the code, its context, corresponding labels, and we construct zero-shot chain-of-thought (CoT) [51] prompts with different command formats for training. In the inference phase, Reasoner outputs multiple explanations based on the majority-voted label from Detector. We constructed two types of prompts: the first type includes the label, code information, and its function call relationships; the second type includes only the label and code information. In §IV, we will investigate the impact of including function call relationships or not. For each type, we designed five different instruction formats for the prompts, totaling ten inference paths, as illustrated in Fig. 1.
Below is an instruction that describes a reasoning task.
### Instruction: [Task Instruction] ### Input: : “‘Solidiy {code} “‘ ### As a Caller: (Optional) “‘ {caller info} “‘ ### As a Callee: (Optional)
“‘ {callee info} “‘ ### Response:
@@
The prompt template used by the Reasoner is shown in Fig. 3. “code”, “caller info”, “callee info”, and “Target Reason” are placeholders for the input code, caller context, callee context, and the target output, respectively. The right table (Reasoner’s Multiple Prompts) in Fig. 4 details the first prompt type with calling context, including [Task Description] denoted as a, [Task Instruction] denoted as b, [Input Description] denoted as c, [Caller Description] denoted as d, [Callee Description] denoted as e, and [Label Information + Zero-shot-CoT Tip] denoted as f. For the second prompt type, the calling context is omitted. Reasoner employed the same fine-tuning method as Detector, as shown by Eq. 1. During inference, the proposed input prompt follows the training format, and Reasoner generates ten answers to interpret Detector’s assessment.
@@ III-D High-quality Training Data Collection and Enhancement
However, to train our model, we also need non-vulnerable benign code (i.e., negative samples), but this type of data is missing in the audit reports. Therefore, we propose our own data enhancement method to derive high-quality negative samples. Specifically, we adopt the GPT-4-based approach described in LLM4Vuln [15] to extract vulnerability knowledge from vulnerability reports on Code4rena. This includes the functionality descriptions of vulnerable functions and the code-level reasons why the vulnerabilities occur. We then cluster this raw vulnerability knowledge based on the functionality descriptions into groups and use GPT-4 to summarize a functionality description for each group. With the hierarchical information of group functionality, individual functionality, and vulnerability negligence, we employ the hierarchical GPT-based matching (i.e., matching the group first, then matching functionality and negligence) in GPTScan [13] to obtain the label information for tested code. A function is labeled as a negative sample if no vulnerability information matches. All prompts used are from LLM4Vuln and GPTScan.
Eventually, we collected a balanced dataset with 1,734 positive samples and 1,810 negative samples. This dataset was divided into training, validation, and test subsets, containing 2,268, 567, and 709 entries, respectively. During training, we use the training and validation sets. During testing, we use the test set.

### v2

# Combining Fine-tuning and LLM-based Agents for Intuitive Smart Contract Auditing with Justifications

@@ Abstract
In this paper, we propose iAudit, a general framework that combines fine-tuning and LLM-based agents for intuitive smart contract auditing with justifications. Specifically, iAudit is inspired by the observation that expert human auditors first perceive what could be wrong and then perform a detailed analysis of the code to identify the cause. As such, iAudit employs a two-stage fine-tuning approach: it first tunes a Detector model to make decisions and then tunes a Reasoner model to generate causes of vulnerabilities. However, fine-tuning alone faces challenges in accurately identifying the optimal cause of a vulnerability. Therefore, we introduce two LLM-based agents, the Ranker and Critic, to iteratively select and debate the most suitable cause of vulnerability based on the output of the fine-tuned Reasoner model. To evaluate iAudit, we collected a balanced dataset with 1,734 positive and 1,810 negative samples to fine-tune iAudit. We then compared it with traditional fine-tuned models (CodeBERT, GraphCodeBERT, CodeT5, and UnixCoder) as well as prompt learning-based LLMs (GPT4, GPT-3.5, and CodeLlama-13b/34b). On a dataset of 263 real smart contract vulnerabilities, iAudit achieves an F1 score of 91.21% and an accuracy of 91.11%. The causes generated by iAudit achieved a consistency of about 38% compared to the ground truth causes.
11footnotetext: Corresponding author: Daoyuan Wu. Work conducted while at NTU.
@@ I Introduction
Fine-tuning [17, 18] could be a promising approach to embed Solidity-specific vulnerability data into the model itself, compared to RAG [19], and thus address the problem mentioned above. In particular, by fine-tuning an LLM with vulnerable and non-vulnerable code, it could effectively perceive whether a new piece of code is vulnerable or not. According to insights from a million-dollar-earning hacker mentioned in the prologue, such intuition is quite important for vulnerability auditing. As such, instead of fine-tuning a single model to generate both vulnerability decisions (i.e., Yes or No) and the causes of vulnerabilities (i.e., the type or reason) simultaneously, we propose a novel two-stage fine-tuning approach. This approach first tunes a Detector model to make decisions only, and then tunes a Reasoner model to generate the causes of vulnerabilities. In this way, the fine-tuned LLMs could mimic human hackers by first making intuitive judgments and then performing follow-up analysis of the code to identify the reasons for vulnerabilities.
@@
We implement this “perception-then-analysis” fune-tuning into a general framework called iAudit 222iAudit is deployed as an auditing module of MetaTrust Labs’ TrustLLM; see https://huggingface.co/MetaTrustSig. for intuitive smart contract auditing. In this implementation, iAudit allows Detector to make multiple intuitive judgments, each representing one perception. To achieve this, iAudit generates multiple variant prompts for the same vulnerability label to tune Detector and similarly employs multiple variant prompts for the same vulnerability reason to tune Reasoner. While it is possible to determine the optimal decision based on majority voting, fine-tuning alone cannot identify the optimal cause for a vulnerability during the inference phase. To address this new problem, we introduce the concept of LLM-based agents to the paradigm of fine-tuning in iAudit. Specifically, we introduce two dedicated LLM-based agents, the Ranker and Critic agents, to iteratively select and debate the most appropriate cause of vulnerability based on the output of the fine-tuned Reasoner model.
@@
To obtain high-quality data for training and testing iAudit, we propose leveraging reputable auditing reports to collect positive samples and employing our own data enhancement method to derive negative samples. Eventually, we collected a balanced dataset consisting of 1,734 positive samples, i.e., vulnerable functions with reasons from 263 smart contract auditing reports, and 1,810 negative samples, i.e., non-vulnerable benign code. We then compared iAudit with traditional full-model fine-tuning methods, including CodeBERT, GraphCodeBERT, CodeT5, and UnixCoder, as well as with prompt learning-based LLMs, such as GPT-4/GPT-3.5 and CodeLlama-13b/34b. Our experimental results show that iAudit achieved an F1 score of 91.21%, significantly outperforming prompt learning-based LLMs (which are in the range of 60%+) and also notably beating other fine-tuned models (which are in the range of 80%+) that used the same training data as ours. Furthermore, in terms of alignment with ground-truth explanations, iAudit’s output is clearly superior to that of other models, reaching a consistency rate of 37.99%. In contrast, the second-ranked GPT-4 achieves only 24%.
@@
Besides the evaluation results, we also conducted three ablation studies to further justify iAudit’s two-stage fine-tuning and majority voting strategies, as well as to measure the impact of additional call graph information on the model’s performance. We summarize the key findings as follows:
@@
  • iAudit’s two-stage approach achieved better detection performance than the integration model, which outputs labels and reasons simultaneously. We also experimentally confirmed that the model struggles to focus on the labels when required to output both types of information.
@@
Roadmap. The rest of this paper is organized as follows. We first introduce the relevant background in Sec. II, followed by the design of iAudit in Sec. III. We then present our experimental setup and the results in Sec. IV. After that, we discuss related work and the limitations in Sec. V and Sec. VI, respectively. Finally, Sec. VIII concludes this paper.
@@ II-C Smart Contracts and Their Vulnerabilities
Figure 1: An overview of iAudit, featuring its four roles: Detector, Reasoner, Ranker, and Critic.
III Design of iAudit
As motivated in Sec. I, iAudit employs a novel two-stage fine-tuning approach and combines it with LLM-based agents for intuitive smart contract auditing with justifications. As shown in Fig. 1, iAudit has the following four roles:
@@ III Design of iAudit
  • Reasoner takes the initial vulnerability perception from Detector to further analyze the potential causes of the vulnerability based on Detector’s decision. By connecting Detector’s output with Reasoner’s reasoning during both training and inference, iAudit achieves two-stage fine-tuning.
  • To identify the optimal cause of a vulnerability during the inference phase, we further introduce the concept of LLM-based agents into the fine-tuning paradigm in iAudit. Specifically, Ranker evaluates the reasons for each potential vulnerability, selecting a top explanation, while Critic further assesses Ranker’s output to debate and determine the most appropriate cause of the vulnerability.
Challenges. While iAudit’s four roles in Fig. 1 are intuitive, training and coordinating them well for effective smart contract auditing with reasonable justifications is difficult. More specifically, we encountered the following four challenges during the design and implementation of iAudit:
@@
How to collect and derive high-quality training data? For a fine-tuned model like iAudit, obtaining high-quality training data is always crucial. We propose leveraging reputable auditing reports to collect positive samples and employing our own data enhancement method to derive negative samples. Since this part is independent of iAudit’s design, we defer its presentation to the end of this section in Sec. III-D.
@@
How to make effective vulnerability judgements? While fine-tuning a model with vulnerable and non-vulnerable code is straightforward, tuning it to be effective with limited data presents a challenge. We make an effort towards addressing this problem in Sec. III-A by opting to use multiple prompts for fine-tuning rather than a single prompt. The advantages of this approach are twofold: (i) it enriches the training dataset by increasing the volume of data, and (ii) it diminishes the bias associated with a single prompt, thereby enhancing the reliability of the results [48]. Optimal vulnerability perception could thus be achieved through majority voting.
@@
How to effectively connect Detector’s vulnerability sensing with Reasoner’s vulnerability reasoning? The fine-tuning of iAudit is unique because it employs a two-stage fine-tuning approach with the Detector and Reasoner models. Therefore, how to effectively connect these two models becomes a new issue not encountered in traditional fine-tuning. We present this aspect of iAudit’s design in Sec. III-B.
@@
How to obtain the optimal vulnerability cause from Reasoner’s output? Since Reasoner also employs multi-prompt fine-tuning, it is necessary to identify the optimal cause of vulnerability among the multiple causes output by Reasoner. We introduce two LLM-based agents, namely the Ranker and Critic components, in Sec. III-C, to iteratively select and debate the most appropriate cause of vulnerability.
An Example of Workflow. To wrap up, Fig. 1 also illustrates an example of iAudit’s workflow. Initially, Detector perceives code vulnerabilities using five different inference paths (prompts). The perceived results are then subjected to majority voting to determine a consensus label. Based on the voting result, Reasoner interprets this outcome according to different inference paths, resulting in ten answers (each considering the context of the code location or not). Next, Ranker selects Reason 1 with a confidence score 9/10 and explains this choice. Critic challenges this choice and advises Ranker to re-evaluate. Taking Critic’s feedback into account, Ranker re-ranks the ten reasons and selects Reason 3 with a confidence score of 10/10. Critic reviews Ranker’s choice again and agrees with this decision. The loop is completed, and the final reason is returned to the user.
@@ III-A Using Multi-prompt Tuning and Majority Voting for Effective Vulnerability Judgements in the Detector
Detector is a fine-tuned expert model responsible for assessing whether code poses any risk. It mimics human intuitive judgment upon seeing a piece of code, assessing whether there are any issues. We employed LoRA [39] to fine-tune CodeLlama-13b [49] in the instruction manner [50] based on a high quality of dataset. During training, for the same input code, we wrap it with multiple prompts. These prompts, with different instruction formats, represent the different inference paths, as illustrated in Fig. 1. In the inference phase of Detector, based on the output results of each prompt, we adopt a majority voting approach to determine the input label and use the voting ratio as the confidence score. Based on Detector’s majority voting result, Reasoner in Sec. III-B then generates different reasons according to different inference paths.
It is worth noting that for the choice of the base model, we randomly selected 16 real logical vulnerabilities to evaluate three popular open-source models: StarCoder, Llama2, and CodeLlama. Upon manual review, we found that StarCoder sometimes refused to respond, and Llama2 provided one contradictory response with two labels. In contrast, CodeLlama offered a more stable response, which led us to choose CodeLlama as the foundational model for further fine-tuning.
Detector’s Prompt Template Below is an instruction that describes a classification task. [Task Description] ### Instruction: [Task Instruction] ### Input: [Input Description]:
“‘Solidiy
{code}
### Response:
@@ III-B Connecting Detector for Reasoner’s Tuning & Inference
Reasoner is an expert model responsible for reasoning about and explaining code vulnerabilities. It interprets the majority voting result of the Detector, generating multiple alternative explanations. During the Reasoner’s LoRA fine-tuning process, our inputs include the code, its context, corresponding labels, and we construct zero-shot chain-of-thought (CoT) [51] prompts with different command formats for training. In the inference phase, Reasoner outputs multiple explanations based on the majority-voted label from Detector. We constructed two types of prompts: the first type includes the label, code information, and its function call relationships; the second type includes only the label and code information. In Sec. IV, we will investigate the impact of including function call relationships or not. For each type, we designed five different instruction formats for the prompts, totaling ten inference paths, as illustrated in Fig. 1.
Reasoner’s Prompt Template Below is an instruction that describes a reasoning task. [Task Description] ### Instruction: [Task Instruction] ### Input: [Input Description]:
“‘Solidiy
{code}
### As a Caller: (Optional)
[Caller Description] “‘
{caller info}
### As a Callee: (Optional)
[Callee Description] “‘
{callee info}
### Response:
[Label Information + Zero-shot-CoT Tip]
{Target Reason}
@@
The prompt template used by the Reasoner is shown in Fig. 3. “code”, “caller info”, “callee info”, and “Target Reason” are placeholders for the input code, caller context, callee context, and the target output, respectively. The right table (Reasoner’s Multiple Prompts) in Fig. 4 details the first prompt type with calling context, including [Task Description] denoted as a, [Task Instruction] denoted as b, [Input Description] denoted as c, [Caller Description] and [Callee Description] denoted as d, and [Label Information + Zero-shot-CoT Tip] denoted as e. For the second prompt type, the calling context is omitted. Reasoner employed the same fine-tuning method as Detector, as shown by Eq. 1. During inference, the proposed input prompt follows the training format, and Reasoner generates ten answers to interpret Detector’s assessment.
@@ III-D High-quality Training Data Collection and Enhancement
However, to train our model, we also need non-vulnerable benign code (i.e., negative samples), but this type of data is missing in the audit reports. Therefore, we propose our own data enhancement method to derive high-quality negative samples. Specifically, we adopt the GPT-4-based approach described in LLM4Vuln [15] to extract vulnerability knowledge from vulnerability reports on Code4rena. This includes the functionality descriptions of vulnerable functions and the code-level reasons why the vulnerabilities occur. We then cluster this raw vulnerability knowledge based on the functionality descriptions into groups using Affinity Propagation [56] as described in [57] and use GPT-4 to summarize a functionality description for each group. With the hierarchical information of group functionality, individual functionality, and vulnerability negligence, we employ the hierarchical GPT-based matching (i.e., matching the group first, then matching functionality and negligence) in GPTScan [13] to obtain the label information for tested code. A function is labeled as a negative sample if no vulnerability information matches. All prompts used are from LLM4Vuln and GPTScan.
Eventually, we collected a balanced dataset with 1,734 positive samples and 1,810 negative samples. In this dataset, vulnerable functions have a median of 49.5 lines of code and a complexity of 18.5, while safe functions have a median of 35.5 lines of code and a complexity of 13.5. The complexity distributions between the two types are somewhat similar but with a slight difference, indicated by a Kullback-Leibler (KL) divergence value of 0.1. This dataset was divided into training, validation, and test subsets, containing 2,268, 567, and 709 entries, respectively. During training, we use the training and validation sets. During testing, we use the test set.
@@
Note that we chose GPT-3.5 mainly because it achieves a good balance between cost and effectiveness—much cheaper than GPT-4, yet comparable in quality for this specific task during our manual comparison. We also added constraints to our prompts during this process to ensure that the enhanced explanations aligned with the actual vulnerabilities.